# clem:todd: A Framework for the Systematic Benchmarking of LLM-Based Task-Oriented Dialogue System Realisations

### v1

@@ A.1 Prompt Templates
Depending on the specific role, the system message and instructions were adapted to emphasize the required behavior (e.g., simulating user intents, extracting slots, managing dialogue state transitions). Following standard prompting approaches (Brown et al., 2020; Wei et al., 2022; Liu et al., 2023a), all prompts were constructed in a zero-shot setting without the use of explicit examples. Prompts were formatted using a system-message followed by a user-message structure.
While the specific content of the prompts varies depending on the role and task, the overall prompt structure remains consistent across all scenarios. Each prompt is composed of the following components:
@@ A.3 Response Format Issues
Some common types of invalid responses include incorrect key-value structures, missing required fields, and multiple function calls when only a single function call was expected. Figure 5 illustrates such invalid responses observed during evaluation.
@@ A.4 Dialogue Quality Evaluation
To further investigate the performance of the user simulator, we conducted a Turing Test to compare generated dialogues against ground-truth dialogues. This evaluation assesses the dialogues produced by Qwen2.5-32B and LLaMA-3.3-70B relative to ground-truth dialogues for the same task, focusing specifically on perceived naturalness.
Model #Human-Like Dialogues
Qwen2.5-32B 19
Llama-3.3-70B 8
Table 12: Results of the Human Turing Test comparing Qwen2.5-32B and LLaMA-3.3-70B dialogues against ground-truth references, based on naturalness judgments across 50 dialogues.
@@
We developed a simple user interface (see Figure 6) that displays two dialogues side by side, allowing annotators to select which dialogue appears more natural. In each comparison, one dialogue is generated by either Qwen2.5-32B or LLaMA-3.3-70B, and the other is the corresponding ground-truth dialogue from the corpus (Budzianowski et al., 2018).
We recruited a student assistant to perform the annotation. The annotator did not have prior experience in dialogue annotation but had general experience with data labeling tasks. Detailed instructions for the Turing Test were embedded within the web interface (see Figure 6) to guide the evaluation process. The annotation task involved evaluating 100 dialogue pairs and took approximately 150 minutes to complete.
The results are summarized in Table 12. For each model, 50 generated dialogues were evaluated against ground-truth dialogues. Out of 50 comparisons, 19 dialogues generated by Qwen2.5-32B were judged to be more natural, while only 8 dialogues generated by LLaMA-3.3-70B were preferred.

### v2

@@ 6 Conclusion
Acknowledgments
The work reported here has been funded by the Bundesministerium für Bildung und Forschung (BMBF, German Federal Ministry of Research), project "COCOBOTS" (01IS21102A) and Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) grant 423217434 (“RECOLAGE”). We thank the anonymous reviewers for their helpful feedback.
@@ A.1 Prompt Templates
Following standard prompting approaches (Brown et al., 2020; Wei et al., 2022; Liu et al., 2023a), all prompts were constructed in a zero-shot setting without the use of explicit examples. Prompts were formatted using a system-message (emphasizing the required behavior) followed by a user-message structure.
While the specific content of the prompts varies depending on the role and task (e.g., simulating user intents, extracting slots, managing dialogue state transitions), the overall prompt structure remains consistent across all scenarios. Each prompt is composed of the following components:
@@ A.3 Response Format Issues
common types of invalid responses include incorrect key-value structures, missing required fields, and multiple function calls when only a single function call was expected. Figure 4 illustrates such invalid responses observed during evaluation.
@@ A.4 Dialogue Quality Evaluation
To further investigate the performance of the user simulator, we conducted a Turing Test to compare generated dialogues against ground-truth dialogues.
@@
We recruited an annotator to perform the annotation. The annotator did not have prior experience in dialogue annotation but had general experience with data labeling tasks. Detailed instructions for the Turing Test were embedded within the web interface (see Figure 5) to guide the evaluation process. The annotation task involved evaluating 100 dialogue pairs and took approximately 150 minutes to complete.
The results are summarized in Table 1 under TT (Turing Test) metric. For each model, 50 generated dialogues were evaluated against ground-truth dialogues. Out of 50 comparisons, 19 dialogues generated by Qwen2.5-32B were judged to be more natural, while only 8 dialogues generated by LLaMA-3.3-70B were preferred.